OpenStudy (anonymous):

Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

help me please

hero (hero):

What a mess this is

OpenStudy (anonymous):

combine like terms

hero (hero):

It's not as simple as that

OpenStudy (anonymous):

i think we have to break up the problem into two problems

hero (hero):

Yes, we do

hero (hero):

start with adding 10 to all sides

OpenStudy (anonymous):

if the system is 2x+5y=-3 -15x-5y=-10 just make a linear combination of the two equations and they's will be eliminated -13x=-13 x=1 then back sub to get y

OpenStudy (anonymous):

yea its 2x + 5y = −3 and −15x − 5y = −10

hero (hero):

Hey eddie, try using separate lines next time.

hero (hero):

Or at least use a semi-colon to separate two different equations.

OpenStudy (anonymous):

yea sorry about that, my fault

OpenStudy (anonymous):

then get y from the original first equation 2x+5y=-3 2(1)+5y=-3 (sub 1 for x) 5y=-5 (subtract 2 from both sides) y=-1 (divide through by 5) your solution is the ordered pair (1,-1)